OpenStudy (anonymous):

7 girls on a bus being driven to school each girl has 7 bags for each bag there are 7 big cats for every big cat there are 7 little cats how many legs are on the bus? o.O

OpenStudy (anonymous):

None?

OpenStudy (anonymous):

My teacher is weird

OpenStudy (anonymous):

well at least its better than his example of derivatives with kegs of beer

OpenStudy (anonymous):

14+28+28=70 because 7*2=14 girls legs, 7*4= 28 big cat legs, 7*4=28 small cat legs

OpenStudy (kamille):

1) 7*2=14 (girls legs) 2)There are 8 cats in every bag (one big one and 7 small ones). So there are 8*4=32 legs in every bag 3)There are 7 bags like these. so 32*7=224 4)14+224=238 legs in total

OpenStudy (kamille):

@dsmith27.ds you are not right, because you have forgotten 7 bags

OpenStudy (anonymous):

Oh i get this. The answer to what the question is asking is 0 legs because a bus doesnt have legs.

OpenStudy (kamille):

@Chef86 Well, I am not a native speaker, but doesn't "on the bus" mean the same as "inside the bus"?